OpenStudy (anonymous):

Explain how to find the solution to this equation: 64^(x – 3) = 4^(2x)

OpenStudy (maheshmeghwal9):

\[64^{x-3}=4^{2x}=>4^{3(x-3)}=4^{2x}\] So \[3(x-3)=2x=>x=9. Ans........\]

OpenStudy (maheshmeghwal9):

gt it?

OpenStudy (anonymous):

Um not 100%. Can you explain please using sentences what you did there...

OpenStudy (maheshmeghwal9):

ya ! u have right to know this:) NOTE: - When 2 exponents with same base related by the sign of equality, are anywhere; just do this: - that exponents will be equal.

OpenStudy (anonymous):

thanks :)

OpenStudy (maheshmeghwal9):

yw:) so is it 100%

OpenStudy (maheshmeghwal9):

i mean did u gt it 200%?

OpenStudy (anonymous):

yes I got it! I was solving these fine but idk why I got stumped on that one

OpenStudy (maheshmeghwal9):

don't worry about maths when u have OS^_^

OpenStudy (maheshmeghwal9):

OpenStudy*